If a person bikes 2.8miles in 10 minutes,how far can he bike in 1.5 hours
Answer by Theo(13342)

10 miles / 60 = 1/6 hours.
he bikes 2.8 miles in 1/6 hours.
1.5 houors = x * 1/6 hours.
x = 6 * 1.5 = 9.
he travels 2.8 * 9 = 25.2 miles in 1.5 hours.

two ways to solve this.

you have to convert hours to minutes or minutes to hours to get the correct answer.

i converted minutes to hours above.
i'll convert hours to minutes down below to show you that you get the same answer if you do the conversions correctly.

1.5 hours * 60 minutes per hour = 1.5 *I 60 = 90 minutes.

ratio method:

2.8 / 10 = x / 90

cross multiply to get 2.8 * 90 = 10 * x

solve for x to get x = 2.8 * 90 / 10 = 2.8 * 9 = 25.2

rate * time = distance method:

if thd person travels 2.8 miles in 10 minutes, then the formula becomes:

rate * 10 = 2.8

solve for rate to get rate = 2.8 / 10 = .28 miles per minute.

at this rate, if he travels for 90 minutes, the equation becomes:

.28 * 90 = distance

solve for distance to get rate = .28 * 90 = 25.2

the ratio method works well when it is a direct ratio.

if it is an inverse ratio, it could get tricky.

i prefer the rate * time = distance method.

it works every time, whether or not the problem winds up being a direct ratio or an inverse ratio.

this problem involved a direct ratio, so it was relatively easy to solve using the ratio method.

how you choose to do it is up to you.

always check your answer to the original problem to make sure you got it right.